# MAVias: Mitigate any Visual Bias

### v1

@@ Abstract
Mitigating biases in computer vision models is an essential step towards the trustworthiness of artificial intelligence models. Existing bias mitigation methods focus on a small set of predefined biases, limiting their applicability in visual datasets where multiple, possibly unknown biases exist. To address this limitation, we introduce MAVias, an open-set bias mitigation approach leveraging foundation models to discover spurious associations between visual attributes and target classes. MAVias first captures a wide variety of visual features in natural language via a foundation image tagging model, and then leverages a large language model to select those visual features defining the target class, resulting in a set of language-coded potential visual biases. We then translate this set of potential biases into vision-language embeddings and introduce an in-processing bias mitigation approach to prevent the model from encoding information related to them. Our experiments on diverse datasets, including CelebA, Waterbirds, ImageNet, and UrbanCars, show that MAVias effectively detects and mitigates a wide range of biases in visual recognition tasks outperforming current state-of-the-art.
Figure 1: MAVias identifies visual biases through foundational models that extract tags representing visual features and assess relevance to the target class. Then, MAVias encodes these features within the vision-language space and integrates them into a bias-aware framework to train a model that is invariant to such visual biases.
@@ 1 Introduction
Computer Vision (CV) progress has been largely driven by Deep Learning (DL) advances [9, 15] and large-scale datasets [6, 25, 21], enabling models to learn complex patterns and visual features with impressive accuracy. However, alongside this progress, concerns have emerged about biases embedded in these models [32, 29, 11, 38, 41, 44] – often stemming from unintended correlations present in the training data [27, 13, 30, 22]. The correlated attributes that are irrelevant, act as “shortcuts” and can significantly impact the model’s reliability and generalization [37, 12]. It is important to note that in the context of this paper, we define as visual bias any characteristic that does not contribute to defining the target class, which we refer to as “irrelevant”. To address this, several methodologies have been developed. Broadly, these fall into two categories: Bias Label-Aware (BLA) and Label-Unaware (BLU) methods. BLA methods leverage the annotations of the attributes introducing the biases to address them [19, 45, 42, 2]. BLU methods focus on extracting bias pseudo-labels in cases of extreme biases where a bias-proxy model (or bias-capturing classifier) can be trained using the task’s target labels that closely align with the bias labels [16, 31, 39, 28].
@@
While effective in certain contexts, both BLA and BLU methods are limited in their applicability when there are multiple, complex, and possibly unknown biases. Common challenging scenarios include:
@@
Unknown biases. In large-scale, general-purpose CV datasets, such as ImageNet, biases may be difficult to identify and largely remain unknown, as they vary widely between different classes and are not prominent enough to allow training of a bias proxy model. For instance, in the ImageNet9 example presented in Tab. 1, a sample labeled as a dog could introduce biases related to the background scene (e.g., armchair, couch, pillow, red), the color of the dog (e.g., black and white), or accessories like a neckband.
@@
Potential biases beyond a predefined set. The CelebA example in Tab. 1, shows that beyond the hair color, additional biases may be present, such as clothing styles (e.g., business suit, tie).
@@
Poor representation of predefined bias. In some cases, biases are reduced to single labels, such as “rural background” in the UrbanCars dataset. However, as shown in Tab. 1, more nuanced descriptors at the instance level (e.g., path, tree, wood, forest, hydrant, red, etc.) provide a richer context for modeling bias.
@@
Existing BLA and BLU methods are not designed to mitigate such biases, leading to models that are biased and/or do not achieve their optimal generalization potential. To address this, Mitigate Any Visual bias (MAVias) introduces a flexible and scalable solution, capable of identifying and mitigating open-set biases in CV datasets. MAVias begins by extracting descriptive tags that capture various visual features, such as general-purpose objects, actions, scenes, and visual attributes. Note that a sufficiently large and comprehensive vocabulary can cover the vast majority of visual features present in CV datasets, enabling the method to be considered as operating in an open-set setting. Then, these tags are processed by a Large Language Model (LLM) to identify which ones are irrelevant to each of the target classes, leading to a rich set of language-encoded visual biases, text descriptions of visual characteristics that are irrelevant to the classification task at hand. We then translate these biases into vision-language embeddings, project them to the main model’s backbone space and then to the classification layer. This projection layer is trained simultaneously alongside the main model, and during training, the output logits are a linear combination of the logits of the main model and those of the projection layer that capture visual biases. This setup allows the main model to be exposed to the biased features – those representing irrelevant information to the target class – in a controlled way that leads to bias-invariant representations. Overall, MAVias provides an effective, end-to-end solution for identifying and mitigating biases in open-set scenarios. We evaluate the proposed method on several datasets involving single-attribute biases (CelebA, Waterbirds) as well as multi-attribute predefined biases (UrbanCars) demonstrating state-of-the-art performance. Furthermore, experiments were conducted on ImageNet9, involving unknown biases, where the suggested approach demonstrates significant gains (from 5.24% to 10.21%) in terms of accuracy compared to existing competitive approaches.
The main contributions of this paper are the following: (i) a framework for identifying instance-specific open-set potential visual biases in CV datasets, (ii) a learning strategy that exploits foundation models to learn bias-invariant representations, and (iii) an extensive evaluation study including comparative and ablation experiments across 4 thematically diverse datasets demonstrating the effectiveness and general applicability of the proposed method, which outperforms the current state-of-the-art.
@@ 2 Related Work
Several methods recently emerged that leverage text (such as captions, keywords, or tags) for bias detection, highlighting the potential of this approach in this domain. For instance, Say My Name (SaMyNa) [4] is an explainability method that tries to discover model biases through text-based pipelines. Similarly, the Bias-to-Text (B2T) [20] and Language-guided (Lg) [50] frameworks discover biases by extracting common keywords from the captions of misclassified images. These methods aim to identify the main source of bias (e.g., hair color in the CelebA example illustrated in Fig. 1) rather than discovering any potentially biased visual characteristic in an open-set setting. Furthermore, OpenBias [8] introduces a framework for detecting biases in text-to-image generative models by leveraging LLMs to propose potential biases from captions. Although operating in a different domain (text-to-image generation), similarly to MAVias  OpenBias acknowledges the need for discovering potential biases in an open-set setting. However, it solely relies on an LLM to generate specific bias lists for a given caption, raising concerns about whether the LLM’s interpretation of biases aligns with human judgment. In contrast, MAVias takes a more structured and systematic approach; it evaluates each descriptive tag by querying the LLM to determine whether it is directly relevant to the target class using a detailed prompt (refer to the Appendix for system prompt details).
Bias mitigation
@@
Recent efforts to mitigate bias in CV have led to the development of various methodologies, including methods with direct access to the labels of attributes introducing bias (i.e., BLA methods) [19, 45, 42, 16, 2, 37, 35, 24] and methods that do not take advantage of such labels but instead rely on deriving pseudo-labels (i.e., BLU methods) [3, 1, 31, 39, 46]. For instance, Learning Not to Learn (LNL) [19] is a BLA method that discourages the model from predicting the attribute introducing bias, while Bias Contrastive-Bias Balance (BC-BB) [16] and FairKL [2] rely on the bias labels to enforce bias-neutral pairwise similarities between the samples using contrastive learning. Some methods have indirect access to the bias labels by utilizing bias-capturing classifiers trained on different datasets offering bias-related information explicitly [39, 40]. Finally, other methods infer pseudo-labels from the biased vanilla model to identify the biases [31, 1, 5]. While these methods have been effective in mitigating biases, they primarily depend on predefined bias labels or pseudo-labels derived from biased models. On the contrary, this work investigates open-set scenarios and introduces a flexible bias mitigation methodology that can discover and handle multiple, diverse biases without requiring a unified bias-labeling system across all the dataset’s samples.
@@ 3.1 Problem Formulation
Let $\mathcal{D}=\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{N}$ be a dataset consisting of $N$ images, where $\mathbf{x}^{(i)}\in\mathcal{X}$ represents an input image, and $y^{(i)}\in\mathcal{Y}$ is the corresponding target label. The goal is to train a DL model $f_{\boldsymbol{\theta}}:\mathcal{X}\to\mathcal{Y}$, parameterized by $\boldsymbol{\theta}$, to predict the target label $y^{(i)}$ given an image $\mathbf{x}^{(i)}$, while mitigating a set of potential biases, $\mathcal{B}^{(i)}$, present in $\mathbf{x}^{(i)}$ that may lead to biased predictions. In this context the term “potential biases” refers to all visual attributes present in $\mathcal{X}$ that are irrelevant to the target class, such as elements of the background of images.
3.2 Methodology
Dataset ImageNet9 Waterbirds UrbanCars CelebA
Target Class dog bird species car type gender
Bias Type unknown background background and objects hair colour
Extracted Tags armchair, black, chair, couch, dog, neckband, pillow, red, sit, white, bamboo, bamboo forest, bird, blue, branch, green, hide, parrot, perch, sit, stand, stem, tree, yellow, car, path, forest, hydrant, lush, park, red, road, sedan, silver, SUV, tree, white, wood, black, business suit, dress, dress shirt, man, stand, stare, suit, tie, wear,
@@ 3.2.1 Language-driven Bias Modeling
First, MAVias utilizes an image tagging model to extract descriptive tags that represent the visual information present in an image. In many general-purpose CV datasets, biases consist of visual information that can be described in text. Formally, for each image $\mathbf{x}^{(i)}$, we derive a set of tags $\mathcal{T}^{(i)}=\{t^{(i)}_{1},t^{(i)}_{2},\dots,t^{(i)}_{m}\}$, where $m$ represents the total number of tags. These capture various visual attributes, including colors, objects, backgrounds, and other features that can either describe the target class or introduce bias.
The next step is to filter out tags that should not influence the decisions of a robust classifier. To achieve this, we leverage an LLM to identify which tags are irrelevant to the target class $y^{(i)}$. We denote this subset of potential biases as $\mathcal{B}^{(i)}\subseteq\mathcal{T}^{(i)}$. These encapsulate visual features that could lead to biased predictions if considered by the model. To ensure the LLM correctly identifies these irrelevant tags, we carefully design the prompt used for this task by providing precise instructions on what tag should be considered relevant (e.g., physical components, defining features, inherent characteristics, etc.) or not (e.g., background details, lighting, textures, other objects, etc.). Details on the prompt formulation process are provided in the Appendix. Table 1 shows several examples of $\mathcal{B}^{(i)}$ for samples belonging to different datasets.
For each set of irrelevant tags $\mathcal{B}^{(i)}$ associated with an image $\mathbf{x}^{(i)}$, we employ a vision-language model to generate a single embedding $\mathbf{e}^{(i)}\in\mathbb{R}^{d}$, where $d$ is the dimension of the embedding. This is produced using the prompt “a photo of $t^{(i)}_{1},t^{(i)}_{2},\dots,t^{(i)}_{k}$”, where $k\leq m$ is the number of irrelevant tags. This aggregate embedding captures the combined information from all irrelevant tags, providing an instance-level representation of the biased features that could affect the model’s behavior. Comparison with alternative embedding approaches is provided in Appendix.
@@ 3.2.2 Bias Mitigation
We define the main model $f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})$, a DL classifier composed of the following: (i) a backbone $f_{\boldsymbol{\theta}_{bb}}(\mathbf{x}^{(i)})$, which extracts feature representations $\mathbf{h}^{(i)}\in\mathbb{R}^{r}$, where $r$ is the feature vector size; (ii) a classification head $f_{\boldsymbol{\theta}_{c}}(\mathbf{x}^{(i)})$ outputting the logits $\mathbf{z}^{(i)}_{\text{main}}=f_{\boldsymbol{\theta}_{c}}(\mathbf{h}^{(i)})\in\mathbb{R}^{p}$, where $p$ is the number of classes. Thus, the overall main model is expressed as:
$f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})=f_{\boldsymbol{\theta}_{c}}(f_{\boldsymbol{\theta}_{bb}}(\mathbf{x}^{(i)})).$ (1)
In parallel, we introduce a projection layer $g_{\boldsymbol{\phi}}$, parameterized by $\boldsymbol{\phi}$, which takes the visual bias embeddings $\mathbf{e}^{(i)}$ as input and outputs embeddings $\mathbf{b}^{(i)}\in\mathbb{R}^{p}$. Note that $g_{\boldsymbol{\phi}}$ is employed to project $\mathbf{e}^{(i)}$ to the feature space of $\mathbf{h}^{(i)}$. Then the corresponding logits are derived through $\mathbf{z}^{(i)}_{\text{tag}}=f_{\boldsymbol{\theta}_{c}}(\mathbf{b}^{(i)})\in\mathbb{R}^{p}$. The final logits $\mathbf{z}^{(i)}$ for each sample are the addition of the main model logits $\mathbf{z}^{(i)}_{\text{main}}$ and the visual bias logits $\mathbf{z}^{(i)}_{\text{tag}}$:
@@
It is worth noting that bias often arises during DL model training because biased attributes in the training data are easier to learn and thus dominate gradient updates [34]. To counteract this, MAVias incorporates visual bias logits into the main model’s logits, ensuring that as the bias in a sample increases, its impact on gradient updates is reduced. The intuition behind this mechanism is that for bias-aligned samples, the value of $\mathbf{z}_{tag}$ is high, effectively reducing the magnitude of $\mathbf{z}_{main}$ and its contribution to the total logits $\mathbf{z}$. This leads to significantly reduced gradients for these samples. This is supported both empirically in Sec. 4.2 and theoretically in the Appendix. In other words, $\mathbf{z}_{tag}$ assists in decoupling the learning of biased features from the actual task at hand, which allows the main model to focus on the relevant features for the target prediction.
@@
Furthermore, since both the main model and the projection layer are trained concurrently, it is essential to ensure the stability of the training process. To achieve this, we introduce a loss function that combines the classification loss with a regularization term. The classification loss ensures that the combined logits $\mathbf{z}^{(i)}$ accurately predict the target label $y^{(i)}$, while the regularization term aligns the magnitude of the main model’s logits $\mathbf{z}^{(i)}_{\text{main}}$ with those of the visual bias logits $\mathbf{z}^{(i)}_{\text{tag}}$. By doing so, we prevent $g_{\boldsymbol{\phi}}$ from dominating or being overshadowed by $h_{\boldsymbol{\theta}}$.
Specifically, the total loss is computed as:
$\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\phi})=\mathcal{L}_{cls}(\mathbf{z}^{(i)},y^{(i)})+\alpha\cdot\mathcal{L}_{reg}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$
where: (i) $\mathcal{L}_{cls}(\mathbf{z}^{(i)},y^{(i)})$ is the classification loss (e.g., cross-entropy loss) between the final logits $\mathbf{z}^{(i)}$ and the ground truth label $y^{(i)}$; (ii) $\mathcal{L}_{reg}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$ is the regularization term for the norm of the logits, calculated as:

### v2

@@ Abstract
Mitigating biases in computer vision models is an essential step towards trustworthy artificial intelligence systems. Existing bias mitigation methods are limited to predefined biases, preventing their use in visual datasets where multiple, possibly unknown biases exist. To address this limitation, we introduce MAVias, an open-set bias mitigation approach that leverages foundation models to discover spurious associations between visual attributes and target classes. MAVias first captures a wide variety of visual features in natural language via a foundation image tagging model, and then leverages a large language model to select visual features that define the target class, resulting in a set of language-coded potential visual biases. It then translates these biases into vision-language embeddings and introduces an in-processing bias mitigation approach to prevent the model from encoding information related to them. Experiments on diverse datasets, including CelebA, Waterbirds, ImageNet, and UrbanCars, show that MAVias effectively detects and mitigates a wide range of biases in visual recognition tasks, outperforming current state-of-the-art.
Figure 1: MAVias identifies instance-level potential visual biases through foundational models that extract tags representing visual features and assess relevance to the target class. Then, MAVias encodes these features within the vision-language space and integrates them into a bias-aware framework to train a model that is invariant to such visual biases.
@@ 1 Introduction
Computer Vision (CV) progress has been largely driven by Deep Learning (DL) advances [dosovitskiy2020image, ho2020denoising] and large-scale datasets [deng2009imagenet, lin2014microsoft, krishna2017visual], enabling models to learn complex patterns and visual features with impressive accuracy. However, alongside this progress, concerns have emerged about biases embedded in these models [ntoutsi2020bias, mehrabi2021survey, fabbrizzi2022survey, sarridis2023towards, sarridis2024facex, wang2022revise] – often stemming from unintended correlations present in the training data [liu2024devil, gupta2024bias, melnychuk2024bounds, li2024be]. The correlated attributes that are irrelevant act as “shortcuts” and can significantly impact the model’s reliability and generalization [ramos2025, sagawa2019distributionally, gupta2018robot]. It is important to note that in the context of this paper, we define as visual bias any characteristic that does not contribute to defining the target class, which we refer to as “irrelevant”. To address this, several methodologies have been developed. Broadly, these fall into two categories: Bias label-Aware (BA) and label-Unaware (BU) methods. BA methods leverage the annotations of attributes introducing the biases to address them [kim2019lnl, wang2020DI, tartaglione2021end, barbano2022fairkl]. BU methods focus on extracting bias pseudo-labels in cases of extreme biases where a bias-proxy model (or bias-capturing classifier) can be trained using the task’s target labels that closely align with the bias labels [hong2021bb, nam2020LfF, sarridis2023flac, ludebiasing].
@@
While effective in certain contexts, both BA and BU methods have limited applicability when there are multiple, complex, and possibly unknown biases. Common challenging scenarios include:
@@
  • Unknown biases: In large general-purpose CV datasets, such as ImageNet, biases are difficult to identify and largely remain unknown, as they vary widely across different classes and are not prominent enough to allow training of bias proxy models. For instance, in the ImageNet9 example of Tab. 1, a sample labeled as a dog could introduce biases related to the background scene (e.g., armchair, couch, pillow, red), the color of the dog (e.g., black and white), or accessories like a neckband.
@@
  • Potential biases beyond a predefined set: The CelebA example in Tab. 1, shows that beyond the hair color, additional biases may be present, such as clothing styles (e.g., business suit, tie).
@@
  • Poor representation of predefined bias: In some cases, biases are reduced to single labels, such as “rural background” in the UrbanCars dataset. However, as shown in Tab. 1, more nuanced descriptors at the instance level (e.g., path, tree, wood, forest, hydrant, red, etc.) provide a richer representation of bias.
Dataset ImageNet9 Waterbirds UrbanCars CelebA
Target Class dog bird species car type hair color
Bias Type unknown background background and objects gender
Extracted Tags armchair, black, chair, couch, dog, neckband, pillow, red, sit, white bamboo, bamboo forest, bird, blue, branch, green, hide, parrot, perch, sit, stand, stem, tree, yellow car, path, forest, hydrant, lush, park, red, road, sedan, silver, SUV, tree, white, wood black, business suit, dress, dress shirt, man, stand, stare, suit, tie, wear
@@
Existing BA and BU methods are not designed to mitigate such biases, leading to models that are biased and/or do not achieve their optimal generalization potential. To address this, Mitigate Any Visual bias (MAVias) introduces a flexible and scalable solution, capable of identifying and mitigating open-set biases in CV datasets. MAVias begins by extracting descriptive tags that capture various visual features, such as general-purpose objects, actions, scenes, and visual attributes. Recent advancements in image tagging [zhang2024recognize] cover large and comprehensive vocabularies (i.e., >4,000>4,000 tags), effectively meeting the open-set requirements of MAVias. Then, these tags are processed by a Large Language Model (LLM) to identify which ones are irrelevant to each of the target classes, leading to a rich set of language-encoded visual biases, text descriptions of visual characteristics that are irrelevant to the classification task at hand. MAVias translates these biases into vision-language embeddings, projects them to the main model’s backbone space and then to the classification layer. This projection layer is trained simultaneously alongside the main model, and during training, the output logits are a linear combination of the logits of the main model and those of the projection layer that captures visual biases. This setup allows the main model to be exposed to the biased features – those representing irrelevant information to the target class – in a controlled way that leads to bias-invariant representations. Overall, MAVias provides an effective, end-to-end solution for identifying and mitigating biases in open-set scenarios. We evaluate the proposed method on several datasets involving single-attribute biases (CelebA, Waterbirds) as well as multi-attribute predefined biases (UrbanCars), demonstrating state-of-the-art performance. Furthermore, experiments were conducted on ImageNet9, involving unknown biases, where the suggested approach demonstrates significant gains (from 5.24% to 10.21%) in terms of accuracy compared to existing competitive approaches.
The main contributions of this paper are the following:
  • A framework for identifying instance-specific open-set potential visual biases in CV datasets.
  • A learning strategy that exploits foundation models to learn bias-invariant representations that force the under-training model to avoid encoding any number of identified potential biases per training sample.
  • An extensive evaluation study including 4 thematically diverse datasets demonstrating the effectiveness and general applicability of MAVias, which outperforms the state-of-the-art.
MAVias implementation is available as part of the VB-Mitigator library [sarridis2025vbmitigator].
@@ 2 Related Work
Several recent methods leverage text (such as captions, keywords, or tags) for bias detection, highlighting the potential of this approach in this domain. For instance, Say My Name (SaMyNa) [ciranni2024say] is an explainability method that tries to discover model biases through text-based pipelines. Similarly, the Bias-to-Text (B2T) [kim2024keywords] and Language-guided (Lg) [zhao2024language] frameworks discover biases by extracting common keywords from the captions of misclassified images. These methods aim to identify the main source of bias rather than discovering potentially biased visual characteristics in an open-set setting. Furthermore, OpenBias [d2024openbias] is a framework for detecting biases in text-to-image generative models by leveraging LLMs to propose potential biases from captions. Although operating in a different domain (text-to-image generation), similarly to MAVias  OpenBias acknowledges the need for discovering biases in an open-set setting. However, it provides an LLM with text-only image descriptions asking for potential types of bias thus missing visual grounding of the depicted information, which could offer essential context and semantics. In contrast, MAVias takes a more structured and systematic approach; it evaluates each descriptive tag by querying the LLM to determine whether it is directly relevant to the target class using a detailed prompt (specified in the supp. material).
Furthermore, bias identification methods typically use a vanilla model and validation data to infer specific biases, while MAVias focuses on defining irrelevant visual features a priori, and leveraging them in model training to mitigate them. While some open-set identification methods can provide bias labels to define subgroups for use by existing bias mitigation approaches [ciranni2024say], it is known that mitigation methods struggle with multi-attribute biases and cannot scale beyond single-attribute subgrouping [li2023whac]. In contrast, MAVias uses instance-level irrelevant visual features without relying on dataset statistics (i.e., neither bias labels nor subgroups w.r.t. them are defined), allowing it to handle complex biases effectively.
Bias mitigation.
@@
Recent bias mitigation methods include those with direct access to the labels of attributes introducing bias (i.e., BA methods) [kim2019lnl, wang2020DI, tartaglione2021end, hong2021bb, barbano2022fairkl, sagawa2019distributionally, qiu2023simple, li2023whac] and those that do not take advantage of such labels but instead rely on deriving pseudo-labels (i.e., BU methods) [cadene2019rubi, bahng2020rebias, nam2020LfF, sarridis2023flac, wu2023discover, yang2024identifying, venkataramani2024causal]. For instance, Learning Not to Learn (LNL) [kim2019lnl] is a BA method that discourages the model from predicting the attribute introducing bias, while Bias Contrastive-Bias Balance (BC-BB) [hong2021bb] and FairKL [barbano2022fairkl] rely on the bias labels to enforce bias-neutral pairwise similarities between the samples using contrastive learning. Some methods have indirect access to the bias labels by utilizing bias-capturing classifiers trained on different datasets offering bias-related information explicitly [sarridis2023flac, sarridis2024badd]. Finally, other methods infer pseudo-labels from the biased vanilla model to identify the biases [nam2020LfF, bahng2020rebias, clark2019LM]. While these methods have been effective in mitigating biases, they primarily depend on predefined bias labels or pseudo-labels derived from biased models. On the contrary, this work explores open-set scenarios and introduces a flexible bias mitigation approach that can discover and handle multiple, diverse biases without requiring a bias-labeling system.
@@ 3.1 Problem Formulation
Let $\mathcal{D}=\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{N}$ be a dataset consisting of $N$ images, where $\mathbf{x}^{(i)}\in\mathcal{X}$ represents an input image, and $y^{(i)}\in\mathcal{Y}$ is the corresponding target label. The goal is to train a DL model $f_{\boldsymbol{\theta}}:\mathcal{X}\to\mathcal{Y}$, parameterized by $\boldsymbol{\theta}$, to predict the target label $y^{(i)}$ given an image $\mathbf{x}^{(i)}$, while mitigating a set of potential biases, $\mathcal{B}^{(i)}$, present in $\mathbf{x}^{(i)}$ that may lead to biased predictions. In this context, the term “potential biases” refers to all visual attributes present in $\mathcal{X}$ that are irrelevant to the target class, such as elements of the background of images or other so-called “shortcuts”.
3.2 Method
@@ 3.2.1 Language-driven Bias Modeling
In many general-purpose CV datasets, biases manifest through visual information that can be described in text. MAVias, first, utilizes an image tagging model to extract tags that describe the visual information present in an image. Formally, for each image $\mathbf{x}^{(i)}$, we derive a set of tags $\mathcal{T}^{(i)}=\{t^{(i)}_{1},t^{(i)}_{2},\dots,t^{(i)}_{m_{i}}\}$, where $m_{i}$ represents the total number of tags for $i$-th sample. These capture various visual attributes, including colors, objects, backgrounds, and other features that can either describe the target class or potentially introduce bias.
The next step is to filter out tags that should not influence the decisions of an unbiased classifier. To achieve this, we leverage an LLM to identify which tags are irrelevant to the target class $y^{(i)}$. We denote this subset of potential biases as $\mathcal{B}^{(i)}\subseteq\mathcal{T}^{(i)}$. These encapsulate visual features that could lead to biased predictions if considered by the model. To ensure the LLM correctly identifies these irrelevant tags, we carefully design the prompt used for this task by providing precise instructions on what tag should be considered relevant (e.g., physical components, defining features, inherent characteristics, etc.) or not (e.g., background details, lighting, textures, other objects, etc.). Details on the prompt formulation process are provided in the supp. material. Table 1 shows several examples of $\mathcal{B}^{(i)}$ for samples belonging to different datasets.
For each set of irrelevant tags $\mathcal{B}^{(i)}$ associated with an image $\mathbf{x}^{(i)}$, we employ a vision-language model to generate a single embedding $\mathbf{e}^{(i)}\in\mathbb{R}^{d}$, where $d$ is the dimension of the embedding. This is produced using the prompt “a photo of $t^{(i)}_{1},t^{(i)}_{2},\dots,t^{(i)}_{k_{i}}$”, where $k_{i}\leq m_{i}$ is the number of irrelevant tags for the $i$-th sample. This aggregate embedding captures the combined information from all irrelevant tags, providing an instance-level representation of the biased features that could affect the model’s behavior. Comparison with alternative embedding approaches is provided in Sec. 4.6.
@@ 3.2.2 Bias Mitigation
We define the main model $f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})$, a DL classifier composed of the following: (i) a backbone $f_{\boldsymbol{\theta}_{bb}}(\mathbf{x}^{(i)})$, which extracts feature representations $\mathbf{h}^{(i)}\in\mathbb{R}^{r}$, where $r$ is the feature vector size; (ii) a classification head $f_{\boldsymbol{\theta}_{c}}(\mathbf{x}^{(i)})$ outputting the logits $\mathbf{z}^{(i)}_{\text{main}}=f_{\boldsymbol{\theta}_{c}}(\mathbf{h}^{(i)})\in\mathbb{R}^{p}$, where $p=\lvert\mathcal{Y}\rvert$. The overall main model is expressed as $f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})=f_{\boldsymbol{\theta}_{c}}(f_{\boldsymbol{\theta}_{bb}}(\mathbf{x}^{(i)}))$. In parallel, we introduce a projection layer $g_{\boldsymbol{\phi}}$, parameterized by $\boldsymbol{\phi}$, which takes the visual bias embeddings $\mathbf{e}^{(i)}$ as input and outputs embeddings $\mathbf{b}^{(i)}\in\mathbb{R}^{p}$. Note that $g_{\boldsymbol{\phi}}$ is employed to project $\mathbf{e}^{(i)}$ to the feature space of $\mathbf{h}^{(i)}$. Then the corresponding logits are derived through $\mathbf{z}^{(i)}_{\text{tag}}=f_{\boldsymbol{\theta}_{c}}(\mathbf{b}^{(i)})\in\mathbb{R}^{p}$. The final logits $\mathbf{z}^{(i)}$ for each sample are the addition of the main model logits $\mathbf{z}^{(i)}_{\text{main}}$ and the visual bias logits $\mathbf{z}^{(i)}_{\text{tag}}$:
@@
It is worth noting that bias often arises during DL model training because biased attributes in the training data are easier to learn and thus dominate gradient updates [pezeshki2021gradient]. To counteract this, MAVias incorporates visual bias logits into the main model’s logits, ensuring that as the bias in a sample increases, its impact on gradient updates is reduced. The intuition behind this mechanism is that for bias-aligned samples, the value of $\mathbf{z}_{tag}$ is high, effectively reducing the magnitude of $\mathbf{z}_{main}$ and its contribution to the total logits $\mathbf{z}$. This leads to significantly reduced gradients for these samples. This is supported both empirically in Sec. 4.5 and theoretically in the supp. material. In other words, $\mathbf{z}_{tag}$ assists in decoupling the learning of biased features from the actual task at hand, which allows the main model to focus on the relevant features for the target prediction.
@@
Furthermore, since both the main model and the projection layer are trained concurrently, it is essential to ensure the stability of the training process. To achieve this, we introduce a loss function that combines the classification loss with a logit alignment term. The classification loss ensures that the combined logits $\mathbf{z}^{(i)}$ accurately predict the target label $y^{(i)}$, while the alignment term controls the relative magnitudes of the main model’s logits $\mathbf{z}^{(i)}_{\text{main}}$ and the visual bias logits $\mathbf{z}^{(i)}_{\text{tag}}$. By doing so, we prevent $g_{\boldsymbol{\phi}}$ from dominating or being overshadowed by $h_{\boldsymbol{\theta}}$. Specifically, the loss is computed as: $\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\phi})=\mathcal{L}_{cls}(\mathbf{z}^{(i)},y^{(i)})+\alpha\cdot\mathcal{L}_{align}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$, where: (i) $\mathcal{L}_{cls}(\mathbf{z}^{(i)},y^{(i)})$ is the classification loss (e.g., cross-entropy loss) between the final logits $\mathbf{z}^{(i)}$ and the ground truth label $y^{(i)}$; (ii) $\mathcal{L}_{align}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$ is the logit alignment term for the norm of the logits, calculated as: